# Constants of motion in gravitational self-force theory

### v1

@@ Abstract
Synergies between self-force theory and other approaches to the gravitational two-body problem have traditionally relied on calculations of gauge-invariant observables as functions of orbital frequencies. However, in self-force theory one can also define a complete set of constants of motion: energy, azimuthal angular momentum, and radial and polar actions. Here we outline how directly utilizing these constants allows for more straightforward comparisons and hybridizations across the parameter space, as well as more streamlined waveform generation through flux-balance laws. Restricting to the case of nonspinning binaries and first order in self-force, we compute the constants of motion and the corrections to fundamental frequencies numerically as well as analytically (to 9PN in a post-Newtonian expansion), establishing consistency with the highest-order (4PN) results available from post-Newtonian theory. We also apply the results to identify the perturbed locations of special curves in the parameter space: circular orbits and the separatrix between bound and plunging orbits.
@@ II.2 Hamiltonian framework and conserved quantities
along with dimensionless frequencies defined via $\Omega_{A}=m_{1}\hat{\Omega}_{A}$. Expanding $\hat{P}_{A}$ and $\hat{J}_{A}$ in powers of $\varepsilon$ at fixed $\hat{\Omega}_{A}$ yields
@@ III.1 Exact expressions
The evolution of the coordinate time, azimuthal phase, and proper time along the orbit can written as
@@ V.1 Circular orbits
Figure 4: Plot of $\delta e_{\odot}^{2}(p)$ using various schemes. The blue dots and error bars are obtained by evaluating Eq. (62) using our numerical results for $\hat{J}_{r(1)}^{\odot}(p)$. The green dot-dashed line is obtained by replacing $\rho(x)$ by its 9PN expression in Eq. (64), whereas the orange dashed line is obtained by using a non-resummed 9PN expansion (71) for $\delta e_{\odot}^{2}(p)$. The vertical grey dotted lines correspond, from left to right, to $p=6$, $p=p_{\text{iso}}\approx 6.38$ and $p=p_{\text{sign}}\approx 6.95$. Thus, for $p_{\text{iso}}<p<p_{\text{sign}}$, we have $\delta e_{\odot}^{2}(p)<0$ and $e_{\odot}$ is an imaginary number.
@@ V.2 Homoclinic orbits on the separatrix
Finally, we note that in an effective-potential description, points on the separatrix correspond to orbits whose energies are at the peak of the effective potential. For example, at geodesic order, all orbits satisfy a radial equation with the familiar form $\left({\mathrm{d}r_{(0)}}/{\mathrm{d}\tau}\right)^{2}=\hat{E}_{(0)}^{2}-V^{(0)}_{\rm eff}(\hat{r},\hat{L}_{(0)})$ (recalling $\hat{r}:=r_{(0)}/m_{1}$). Unstable circular orbits lie precisely at the peak of the potential, with $\hat{E}^{2}_{(0)}=V^{(0)}_{\rm eff}(\hat{r}_{\rm min}(\hat{L}_{(0)}),\hat{L}_{(0)})$, and the homoclinic orbits are eccentric trajectories with this same $\hat{E}_{(0)}$ and $\hat{L}_{(0)}$. We can then describe the 0SF separatrix by the curve $\hat{E}^{\star}_{(0)}(\hat{L}_{(0)})$, as given by the 0SF relations in Eq. (78). If we neglect dissipation, then this picture persists at 1SF order, with a perturbed potential $V_{\rm eff}(r_{p}/m_{1},\hat{L},\varepsilon)$. We can then describe the energy of orbits on the separatrix as the energy at the peak of the perturbed effective potential, which we write as
@@ V.3 Innermost stable circular orbit
We now specialize to the circular case and verify that we are able to recover known results for the shift in the frequency of the innermost stable circular orbit (ISCO), as well as the values of constants of motion at the ISCO. Most of these quantities are highly singular near the ISCO: limits do not necessarily commute, Taylor-expansions are often singular when expanding around $(p,e)=(6,0)$ and different orders of operations can yield different results. Thus, we here follows the usual definitions in the literature: we first restrict to the curve of circular orbits, then take the limit towards the ISCO.
@@
Decomposing the reduced ISCO frequency in powers of the mass ratio, namely as $\hat{\Omega}_{\text{ISCO}}=\hat{\Omega}_{(0)}^{\text{ISCO}}+\varepsilon\,\hat{\Omega}_{(1)}^{\text{ISCO}}+\mathcal{O}(\varepsilon^{2})$, we then find that the relative ISCO frequency shift reads
@@
After properly accounting for the different conventions, we find very good agreement with the highly accurate redshift for circular orbits used in the 1PA waveform model [167]; good agreement in also found with Refs. [99, 104]

### v2

@@ Abstract
Synergies between self-force theory and other approaches to the gravitational two-body problem have traditionally relied on calculations of gauge-invariant observables as functions of orbital frequencies. However, in self-force theory one can also define a complete set of constants of motion: energy, azimuthal angular momentum, and radial and polar actions. Here we outline how directly utilizing these constants allows for more straightforward comparisons and hybridizations across the parameter space, as well as more streamlined waveform generation through flux-balance laws. Restricting to the case of eccentric, nonspinning binaries and first order in self-force, we compute the constants of motion and the corrections to fundamental frequencies numerically as well as analytically (to 9PN in a post-Newtonian expansion), establishing consistency with the highest-order (4PN) results available from post-Newtonian theory. We also apply the results to identify the perturbed locations of special curves in the parameter space: circular orbits and the separatrix between bound and plunging orbits.
@@ II.2 Hamiltonian framework and conserved quantities
along with dimensionless frequencies $\hat{\Omega}_{A}=m_{1}\Omega_{A}$. Expanding $\hat{P}_{A}$ and $\hat{J}_{A}$ in powers of $\varepsilon$ at fixed $\hat{\Omega}_{A}$ yields
@@ III.1 Exact expressions
The evolution of the coordinate time, azimuthal phase, and proper time along the orbit can be written as
@@ V.1 Circular orbits
Figure 4: Plot of $\delta e_{\odot}^{2}(p)$ using various schemes. The blue dots and error bars are obtained by evaluating Eq. (62) using our numerical results for $\hat{J}_{r(1)}^{\odot}(p)$. The green dot-dashed line is obtained by replacing $\rho(x)$ by its 9PN expression in Eq. (64), whereas the orange dashed line is obtained by using a non-resummed 9PN expansion (71) for $\delta e_{\odot}^{2}(p)$. The vertical grey dotted lines correspond to, from left to right, $p=p_{\text{iso}}\approx 6.38$ and $p=p_{\text{sign}}\approx 6.95$. Thus, for $p_{\text{iso}}<p<p_{\text{sign}}$, we have $\delta e_{\odot}^{2}(p)<0$ and $e_{\odot}$ is an imaginary number.
@@ V.2 Homoclinic orbits on the separatrix
Finally, we note that in an effective-potential description, points on the separatrix correspond to orbits whose energies are at the peak of the effective potential. That textbook description of the geodesic dynamics extends straightforwardly to 1SF order: The condition $g^{\alpha\beta}p_{\alpha}p_{\beta}=(m_{2})^{2}$, together with $p_{t}=E+\varepsilon\delta p_{t}$ and $p_{\phi}=L+\varepsilon\delta p_{\phi}$, implies the (adimensionalized) radial equation
$\left(\frac{d\hat{r}_{p}}{d\hat{\tau}}\right)^{\!2}=\hat{E}^{2}-V(\hat{r}_{p},\hat{E},\hat{L},\varepsilon),$ (80)
with the effective potential
$V=f_{p}\left(1+\frac{\hat{L}^{2}}{\hat{r}_{p}^{2}}\right)+2\varepsilon\left(f_{p}\frac{\hat{L}\delta\hat{p}_{\phi}}{\hat{r}_{p}^{2}}-\hat{E}\delta\hat{p}_{t}\right).$ (81)
[…]
Just as in the usual geodesic description, Eq. (80) implies
$\hat{E}^{2}=V(\hat{r}_{\rm min/max},\hat{E},\hat{L},\varepsilon)$ (82)
at turning points and on circular orbits. Differentiating Eq. (80) with respect to $\hat{\tau}$ and restricting to the conservative sector (where $E$ and $L$ are constants), we also get
$\frac{d^{2}\hat{r}_{p}}{d\hat{\tau}^{2}}=-\frac{1}{2}\frac{\partial V}{\partial\hat{r}_{p}},$ (83)
again just as in the geodesic case. Unstable circular orbits correspond to maxima of the effective potential, where $\frac{\partial V}{\partial\hat{r}_{p}}=0$ picks out circular orbits and $\frac{\partial^{2}V}{\partial r_{p}^{2}}<0$ ensures they are unstable. The two relations (82) and $\frac{\partial V}{\partial\hat{r}_{p}}=0$, with $\frac{\partial^{2}V}{\partial r_{p}^{2}}<0$, determine $\hat{E}_{\rm UCO}(\hat{L},\varepsilon)$ and $\hat{r}_{\rm UCO}(\hat{L},\varepsilon)$. The homoclinic orbits are eccentric trajectories with $\hat{E}^{\star}(\hat{L},\varepsilon)=\hat{E}_{\rm UCO}(\hat{L},\varepsilon)$ and $\hat{r}_{\rm min}(\hat{L},\varepsilon)=\hat{r}_{\rm UCO}(\hat{L},\varepsilon)$, such that they precisely reach the peak of the effective potential at their periapsis, in the infinite past and future.
[…]
@@ V.3 Innermost stable circular orbit
We now specialize to the circular case and verify that we are able to recover known results for the shift in the frequency of the innermost stable circular orbit (ISCO), as well as the values of constants of motion at the ISCO. Most of these quantities are highly singular near the ISCO: limits do not necessarily commute, Taylor-expansions are often singular when expanding around $(p,e)=(6,0)$ and different orders of operations can yield different results. Thus, we here follow the usual definition in the literature: we first restrict to the curve of circular orbits, then take the limit towards the ISCO. Precisely at the ISCO, the curve of circular orbits (defined by $J_{r}=0$) intersects the separatrix (defined by $\Omega_{r}=0$), and we can equivalently define the ISCO by $\Omega_{r}=0=J_{r}$.
@@
Decomposing the reduced ISCO frequency in powers of the mass ratio, namely as
$\hat{\Omega}_{\text{ISCO}}=\hat{\Omega}_{(0)}^{\text{ISCO}}+\varepsilon\,\hat{\Omega}_{(1)}^{\text{ISCO}}+\mathcal{O}(\varepsilon^{2}),$ (91)
we then find that the relative ISCO frequency shift reads
@@
After properly accounting for the different conventions, we find very good agreement with the highly accurate redshift for circular orbits used in the 1PA waveform model [167]; good agreement is also found with Refs. [99, 104]